What is the pH of an aqueous solution that is 0.011 M HF (Ka = 3.5 × 10−4 ) and 0.015 M NaF?

1. 3.33

2. 3.59

3. 3.46

4. 1.95

5. 5.27


1
Expert's answer
2021-03-12T05:59:34-0500

According to Henderson-Hasselbalch equation:

pH=pKa+logC(salt)/C(acid)

pKa=-logKa=-log3.5*10-4=3.456

pH=3.456+log0.015/0.011=3.456+0.135=3.59.

Answer: 2 - 3.59.